6. Mike Jimenez is president of the teachers' union he would like to investigate the salary structure of classroom teachers in the district. that affect a teacher's salary: the teacher has a master's degree. A random sample of 20 teachers resulted in the following data for Preston School District. In preparing for upcoming negotiations, He believes there are three factors years of experience, a rating of teaching effectiveness given by the principal, and whether Years ofPrincipal's Master's Salary ($ thousands) experience Rati 35 43 S1 60 73 80 52.1 43.6 40.3 53 51.6 57 46.8 58.6 42.8 36.7 40.6 61.8 54.7 48.4 53.6 51.8 47.7 42.8 53.8 54 19 90 30 23 17 12 14 76 25 90 62 80 72 1 -yes, 0 no What is the question being asked? a. Label your dependent variable: b. Label all of the independent variables: C.
d. Develop a correlation matrix. Which independent variable has the strongest correlation with the dependent variable? State the variable and the correlation coefficient. d. Determine the regression equation e. e. What salary would you estimate for a teacher with five years' experience, a rating by the principal of 60 master's degree? f. 0.05 Run a diagnosis to see if the model is significant at α State the null and alternate hypotheses &. Ho: Compute the value of the test statistic h. Compute the p value i. What is your decision regarding the null hypothesis? Interpret the result.
0.05 k. Run a diagnosis to determine if the slope of the individual variables are significant at a P-value Significant? (Yes, No) Variable j. Is Multicollinearity a problem? Compute the VIF for each independent variable. VIF Value Problem? (Yes, No) Variable k. Overall Conclusions:
Answer #1

a). Dependency of Salary on Years of experience, Principal's rating and Master's Degree.

b). Salary is the dependent variable

c). Years of experience, Principal's rating and Master's Degree are the independent variable

d).

Salary Years of experience Principal's rating Master's Degree
Salary 1.000
Years of experience 0.900 1.000
Principal's rating 0.335 0.204 1.000
Master's Degree 0.857 0.725 0.207 1.000

Years of experience has the strongest correlation with Salary i.e 0.9

e). Regression equation is Salary = 37.918+0.623*Years of experience+0.045*Principal's rating+5.443*Master's Degree

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.957
R Square 0.916
Adjusted R Square 0.900
Standard Error 2.118
Observations 20
ANOVA
df SS MS F Significance F
Regression 3 779.918 259.973 57.975 0.000
Residual 16 71.748 4.484
Total 19 851.666
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 37.918 1.669 22.718 0.000 34.380 41.457 34.380 41.457
Years of experience 0.623 0.115 5.421 0.000 0.379 0.866 0.379 0.866
Principal's rating 0.045 0.025 1.784 0.093 -0.008 0.099 -0.008 0.099
Master's Degree 5.443 1.386 3.926 0.001 2.504 8.382 2.504 8.382

f). Years of experience = 5

Principal's Rating = 60

Master's Degree = 0 (No)

Then predicted salary = 37.918+0.623*5+0.045*60+5.443*0 = 43.7 (rounded to 1 decimal place)